a solution containing 0.275 g of NaOH require 35.5ml .of HCl for neutralization. what is the normality of HCl?
1
Expert's answer
2015-11-24T14:53:20-0500
Solution
Neutralization NaOH occurs according to the equation :
NaOH + HCl = NaCl + H2O
From the equation it is seen that 1 mol NaOH is reacted with one mole of HCl.
The molar mass of NaOH is 40.0 g / mol .
Then the sample solution contained
0.275 / 40.0 = 0.006875 mol NaOH.
For neutralization require 0.006875 mol hydrochloric acid . This amount of acid contained in the solution was 35.5 ml . Then the molar concentration of HCl solution is
(0,006875 ∙ 1000) / 35.5 = 0.1937 mol/l
Since HCl - monobasic acid , its normality is molarity .
Normality of HCl is 0.1937 mol - eq/l.

Answer: Normality of HCl is 0.1937 mol - eq/l